# Voyager 1 Data Reveals Signatures of the Local Gas and Cosmic-Ray Source Distributions

### v1

@@ Abstract
We investigate the effects of the nearby interstellar medium (ISM) on the locally measured cosmic‑ray (CR) spectra. Using the GalProp code we explore how variations in the local gas and source distributions affect spectral features at low energies. Comparing with recent Voyager 1 data taken in the local ISM, we show that for a realistic interstellar gas distribution, the nearest source of the low energy CR particles observed nearby the Solar system is constrained to be within the range $\sim$150–200 pc distant. We find that the modelling supports the conclusion of Cummings_2025 that there is a significant fraction of primary Boron in its observed spectrum at low energies. Our study shows that detailed modelling of the immediate Galactic environment is required to robustly infer Galactic CR propagation parameters from local measurements, and that accounting for nearby ISM structure can alleviate tensions between direct CR data and global propagation models.
1 Introduction
Cosmic rays (CRs) are a uniquely sensitive probe of the local interstellar medium (ISM) conditions. The combined influence of nearby CR sources, small-scale variations in the interstellar gas, and the dynamical history of the Solar neighbourhood are encoded in their spectra at both low ($\lesssim$1 GeV) (e.g., 2016Sci...352..677B; 2021ApJ...913....5B, and references therein) and very high ($\gtrsim$1 TeV) energies (e.g., 2024AdSpR..74.4264M, and references therein). For decades, attempts to access the low-energy signatures have been hampered by the modulation of CR spectra by the Solar wind. But measurements by Voyager 1 (V1) beyond the heliopause now offer the first direct view of the local interstellar spectrum down to a few MeV nucleon-1 energies (Cummings_2025). These data open a window onto CR transport on scales of tens to hundreds of parsecs, which are at the scales where the structures of the very local ISM (VLISM) are known to vary considerably compared to averaging over the broader Galactic disc.
A defining feature of the nearby Galactic environment is the Local Bubble (LB): a cavity of hot, tenuous plasma $\sim$50–200 pc around the Solar system, bounded by a complex shell of neutral and molecular material. Recent 3D reconstructions of dust and gas reveal that this cavity is highly irregular (e.g., 2024ApJ...973..136O; 2025A&A...695A. [^2] S) and was likely formed by a sequence of supernova (SN) explosions over the past 10–20 Myr (2016Natur.532...73B). The LB’s very low gas density modifies the dominant loss processes for low-energy CRs, while its apparent lack of recent SN activity implies a deficit of local CR injection. These are coupled consequences of the same evolutionary history, which have so far not been treated together in a self-consistent modelling framework.
@@ 2 GalProp Framework
GalProp has 27+ years of development behind it (1998ApJ...493..694M; 1998ApJ...509..212S). It is consistently updated to keep up with the ever increasing amount and precision of experimental data. The latest release of GalProp, version 57.1, became public in 2022 (2022ApJS..262...30P, and references therein). The dedicated website [^1] https://galprop.stanford.edu provides source releases, supporting data products and run configurations for reproducing published results, and a facility to run GalProp releases via web browsers (2011CoPhC.182.1156V).
The GalProp code solves a system of about 90+ time-dependent transport equations (partial differential equations in 3D or 4D: spatial variables plus energy) with a given source distribution and boundary conditions to give the intensity distributions for all CR species through the ISM: 1H-64Ni, $\bar{p}$, $e^{\pm}$. The spatial boundary conditions assume free particle escape. The propagation equations include terms for convection, distributed reacceleration, energy losses, nuclear fragmentation, radioactive decay, and production of secondary particles and isotopes (for details of these processes and formalism, see 1998ApJ...509..212S).
For a given halo size, the diffusion coefficient $D_{xx}(\rho)$, as a function of rigidity $\rho$, and other propagation parameters can be determined from secondary-to-primary nuclei ratios, typically B/C, [Sc+Ti+V]/Fe, and/or $\bar{p}/p$. If reacceleration is included, the momentum-space diffusion coefficient $D_{pp}$ is related to the spatial coefficient $D_{xx}=\beta D_{0}\rho^{\delta}$ (1994ApJ...431..705S), where $\beta=v/c$ is the particle velocity, $\rho$ is the magnetic rigidity, and $\delta=1/3$ for a Kolmogorov spectrum of interstellar turbulence (1941DoSSR..30..301K), or $\delta=1/2$ for an Iroshnikov-Kraichnan cascade (1964SvA.....7..566I; 1965PhFl....8.1385K), but can also be arbitrary. The spatial diffusion coefficient can also depend on position. This option was developed and utilised by 2015ApJ...799...86A, where the positional dependence of the spatial diffusion coefficient was linked to the distribution of the Galactic magnetic field (GMF) strength.
@@
The GalProp code computes a complete network of primary, secondary, and tertiary isotope production starting from input CR source abundances. The nuclear reaction network is built using the 2018NDS...151D...3., where a detailed description of its method of construction is given by 2020ApJS..250...27B; BoDe21. Included are multistage chains of $p$, $n$, $d$, $t$, 3He, $\alpha$, and $\beta^{\pm}$-decays, and K-electron capture, as well as, in several cases, more complicated reactions. This accounts for up to 81 daughter nuclei in the final state for each fragment produced in the spallation of the target nucleus, plus an unlimited number of $p$, $n$, and $\beta^{\pm}$-decays.
Because the decay branching ratios and half-lives of fully stripped and hydrogen-like ions may differ (a well-known example is 7Be), GalProp includes the processes of K-electron capture, electron pickup from the neutral ISM gas, and formation of hydrogen-like ions as well as the inverse process of electron stripping (1973RvMP...45..273P; 1978PhDT........12W; 1979PhDT........67C). Meanwhile, the fully stripped and hydrogen-like ions are treated as separate species.
The isotopic production cross section routines are built using LANL nuclear codes (2001ICRC....5.1836M; 2003ICRC....4.1969M; 2004AdSpR..34.1288M; 2005AIPC..769.1612M; 1996PhRvC..54.1341B; 1998PhRvC..57..233B), databases (LANL, EXFOR [^2] https://www-nds.iaea.org/exfor/), an extensive literature search (BoDe21), and parameterizations (2003ApJS..144..153W; 1998ApJ...501..911S; 1998ApJ...501..920T; 1983ApJS...51..271L). Cross sections for production of 2H, 3H, 3He, and total inelastic cross sections parameterizations (1996PhRvC..54.1329W; BarPol1994) and corrected Tripathi formalism are described in 2022ApJS..262...30P. The latest updates of the production cross sections are summarised in Cummings_2025.
@@
For 2D calculations, analytical models for the gas density distribution are available (2002ApJ...565..280M) (GP2D). The radial distribution for H i is taken from 1976ApJ...208..346G while the vertical distribution is from 1990ARA&A..28..215D for galactocentric radial distances $0\leq R\leq 8$ kpc and 1986A&A...155..380C for $R\geq 10$ kpc with linear interpolation in between. The CO gas distribution is taken from 1988ApJ...324..248B for 1.5 kpc<<RR<<10 kpc, and from 1990A&A...230...21W for $R\geq$10 kpc, and is augmented with the 2007A&A...467..611F model for $R\leq$1.5 kpc. Both of the 2D atomic and molecular gas distributions are rescaled to the common IAU recommended Sun-Galactic centre (GC) distance of $R_{\odot}=8.5$ kpc (1986MNRAS.221.1023K). The H ii gas distribution is given by the NE2001 model (2002astro.ph..7156C; 2003astro.ph..1598C; 2004ASPC..317..211C) with the updates from 2008PASA...25..184G.
For 3D simulations, the H i and 12CO distributions from 2018ApJ...856...45J (J18) are available. These were developed using a maximum-likelihood forward-folding optimisation applied to the LAB-H i (2005A&A...440..775K) and CfA composite CO data (2001ApJ...547..792D; 2004ASPC..317...66D). New with this paper we also include the gas distribution from 2025A&A...693A.139S (S25) who made a reconstruction of a very similar collection of data to the J18 work using 3D Gaussian processes to model correlations in the interstellar gas over different sight lines to enforce coherent structure for its distribution. Compared to the 2D models, the added degrees of freedom allow the optimised distributions to better reproduce the features observed in the line-emission surveys.
Multiple 2D and 3D models for the interstellar radiation field and GMF are also available with GalProp. However, these affect the propagation/energy losses for the CR electrons/positrons, which are not investigated in the current work. We therefore refer the interested reader to 2022ApJS..262...30P and associated references for more details.
3 Calculations
3.1 Model Configurations
Figure 1: Model configuration components. Left panel: profiles for the CR source and GP2D gas surface densities along the $Y=0$ kpc direction. Line styles: source density, solid; H i surface density, dash-dotted; H2 surface density, long dashed. Right panel: S25 model gas surface density for the nearby Galactic neighbourhood. The Solar system is marked by the Sun symbol in the centre, and the dashed line is a circle of radius 200 pc about it.
@@ 3.1 Model Configurations
The baseline model is 2D Galactocentric axisymmetric with a spatial model for the CR sources following the distribution of pulsars from 2006MNRAS.372.. [^7] L. The injection spectrum for the sources is a power-law in rigidity with a single injection index for all species below a break at 3 GV and a different common index for all species above that. The interstellar gas distribution is the standard 2D model GP2D described above. The propagation model uses a single power-law index $\delta$ for the diffusion coefficient $\propto D_{0}\rho^{\delta}$, where $D_{0}$ is the normalisation at 4 GV and $\rho$ is rigidity. The size of the Galaxy has maximum radius of 20 kpc with a $Z$-boundary (halo size) 4 kpc.
For the 3D models, the PD phenomenology and rigidity dependence for the diffusion coefficient is the same as the baseline. The size of the Galaxy has a maximum $X/Y$ boundary at $\pm$20 kpc and $Z$-boundary at 4 kpc. For the gas spatial distribution we use the GP2D and S25 models described above. The molecular gas density for the S25 is obtained using the $X_{CO}$ distribution used for the baseline/GP2D model. The source density model is the same as the baseline. For the 3D solutions we also consider a spherical hole in the source density about the Solar system. With these models we test the effect of smooth and inhomogeneous source and gas distributions for the nearby Galactic neighbourhood.
@@
All model configuration files necessary to reproduce our results will be available as supplementary material, as well as from the GalProp website as is our usual practice.
3.2 CR Solution Geometry
Figure 2: Model combination results together with V1 data for 2013/1–2021/365 in the VLISM from Cummings_2025 for elemental spectra Z=1–10. Line styles: 2D, 3D/0, solid; 3D/100, short dashed; 3D/200, long dashed; S25/0, dotted; S25/100, short dash dotted; S25/200, long dash dotted.
@@ 3.2 CR Solution Geometry
For 3D models, we calculate the CR solution using a spatial grid with the non-uniform tan grid functions for all coordinates $X,Y,Z$, as described by Appendix E of 2022ApJS..262...30P. The parameters for the individual coordinate functions are chosen so that the pixel size close to the Solar system is $\sim$12 pc. For the $X/Y$ coordinates the parameters are chosen so that the resolution slowly decreases with distance from the Solar system until it reaches the size of the $Z$ (halo) boundary, and then coarsens more beyond that to the respective planar boundaries. Because the gas distribution falls rapidly with increasing $Z$, the grid function parameters are chosen so that the resolution slowly decreases only until a distance 1 kpc above the plane and then coarsens more beyond that to the halo boundary.
@@
3.3 Tuning Procedure
The 2D baseline model was obtained by Cummings_2025 using the GalFit package. This is a tool that iteratively optimises a GalProp model configuration parameters against a user-specfied collection of data. For the 2D case this is computationally feasible because the likelihood evaluations are very quick, allowing for the parameter space to be adequately sampled. Such a procedure is not feasible for the high-resolution 3D models in this paper because the computational costs are much higher.
@@ 3.3 Tuning Procedure
3.4 Results
Figure 3: Model combination results together with V1 data for 2013/1–2021/365 in the VLISM from Cummings_2025 for B/C. Line styles as for Fig. 2.
Figure 4: Model combination results together with V1 data for 2013/1–2021/365 in the VLISM from Cummings_2025 for elemental spectra Z=11–19 (left) and Z=20–28 (right). Line styles as for Fig. 2.
Figure 5: Elemental ratios for He, O, and Fe compared with V1 data. For each ratio, the corresponding V1 ratio has been derived from the spectra using power-law interpolation at the geometric mean energy for the elemental denominator species. Uncertainties are obtained by combining in quadrature the 1-sigma uncertainties for each elemental species interpolated using the same method. Left panel: ratios for combinations with 3D/GP2D and S25 gas models with source density hole sizes 0, 100, and 200 pc. Right panel: ratios for S25 gas model with source density hole sizes 0, 100, 150, 200, and 400 pc. Line styles are same as for Fig. 2, with the addition of double-dot dashed and triple-dotted dashed lines for 150 pc and 400 pc source density holes, respectively.
Figure 6: Model results for baseline and S25/150 with Boron pure secondary (4022a) and with primary Boron following 4022c abundances obtained by Cummings_2025 together with V1 data for 2013/1–2021/365 in the VLISM. Panels: B/C, left; Be/O, right. The V1 Be/O ratio is derived using power-law interpolation at the geometric mean energy for the elemental Be spectrum. Line styles as for Fig. 5, with orange lines showing 4022c model curves.
Figure 2 shows the results for the elemental spectra for Z=1–10 with the V1 data from Cummings_2025. The different line styles correspond to the GP2D and S25 gas models with different source density hole radial sizes: no hole, 100 pc, and 200 pc. It can be seen that the modelled CR spectra for all configurations closely agree for energies $\gtrsim$100–200 MeV nucleon-1 showing consistent normalisation. Differences occur for energies below this according to the gas model used and source density hole size.
Between the gas models, the CR intensities for the homogeneous (no hole) source densities are different. While the GP2D/3D combination is as for the 2D baseline fit, for the S25/0 and S25/100 configurations the low-energy spectra are considerably higher. The CR injection occurs right up to the immediate vicinity of the Solar system when there is no hole. The inhomogeneous gas distribution is particularly low $\lesssim$100 pc nearby, and this results in slower energy losses. Consequently, the spectra are not as attenuated as for the homogeneous gas distribution with the GP2D model.

### v2

@@ Abstract
We investigate the effects of the nearby interstellar medium (ISM) on the locally measured cosmic‑ray (CR) spectra. Using the GalProp code we explore how variations in the local gas and source distributions affect spectral features at low energies. Comparing with recent Voyager 1 measurements taken in the local ISM, we show that for a realistic interstellar gas distribution the data favour models in which there are no significant CR sources within $\sim$150-200 pc of the Solar system, implying that the nearest dominant contributors to the low-energy CR flux are located at distances beyond this range. We find that the modelling supports the conclusion of Cummings et al. (2025) that there is a significant fraction of primary Boron in its observed spectrum at low energies. Our study shows that detailed modelling of the immediate Galactic environment is required to robustly infer Galactic CR propagation parameters from local measurements, and that accounting for nearby ISM structure can alleviate tensions between direct CR data and global propagation models.
I Introduction
Cosmic rays (CRs) are a uniquely sensitive probe of the local interstellar medium (ISM) conditions. The combined influence of nearby CR sources, small-scale variations in the interstellar gas, and the dynamical history of the Solar neighbourhood are encoded in their spectra at both low ($\lesssim$1 GeV) (e.g., Binns et al. 2016; Boschini et al. 2021b, and references therein) and very high ($\gtrsim$1 TeV) energies (e.g., Malkov et al. 2024, and references therein). For decades, attempts to access the low-energy signatures have been hampered by the modulation of CR spectra by the Solar wind. But measurements by Voyager 1 (V1) beyond the heliopause now offer the first direct view of the local interstellar spectrum down to a few MeV nucleon-1 energies (Cummings et al. 2025). These data open a window onto CR transport on scales of tens to hundreds of parsecs, which are at the scales where the structures of the very local ISM (VLISM) are known to vary considerably compared to averaging over the broader Galactic disc.
A defining feature of the nearby Galactic environment is the Local Bubble (LB): a cavity of hot, tenuous plasma $\sim$50–200 pc around the Solar system, bounded by a complex shell of neutral and molecular material. Recent 3D reconstructions of dust and gas reveal that this cavity is highly irregular (e.g., O’Neill et al. 2024; Soler et al. 2025) and was likely formed by a sequence of supernova (SN) explosions over the past 10–20 Myr (Breitschwerdt et al. 2016; Zucker et al. 2022). The LB’s very low gas density modifies the dominant loss processes for low-energy CRs, while its apparent lack of recent SN activity implies a deficit of local CR injection. These are coupled consequences of the same evolutionary history, which have so far not been treated together in a self-consistent modelling framework.
@@ II GalProp Framework
GalProp has 27+ years of development behind it (Moskalenko and Strong 1998; Strong and Moskalenko 1998). It is consistently updated to keep up with the ever increasing amount and precision of experimental data. The latest release of GalProp, version 57.1, became public in 2022 (Porter et al. 2022, and references therein). The dedicated website [^1] https://galprop.stanford.edu provides source releases, supporting data products and run configurations for reproducing published results, and a facility to run GalProp releases via web browsers (Vladimirov et al. 2011).
The GalProp code solves a system of about 90+ time-dependent transport equations (partial differential equations in 3D or 4D: spatial variables plus energy) with a given source distribution and boundary conditions to give the intensity distributions for all CR species through the ISM: 1H-64Ni, $\bar{p}$, $e^{\pm}$. The spatial boundary conditions assume free particle escape. The propagation equations include terms for convection, distributed reacceleration, energy losses, nuclear fragmentation, radioactive decay, and production of secondary particles and isotopes (for details of these processes and formalism, see Strong and Moskalenko 1998).
For a given halo size, the diffusion coefficient $D_{xx}(\rho)$, as a function of rigidity $\rho$, and other propagation parameters can be determined from secondary-to-primary nuclei ratios, typically B/C, [Sc+Ti+V]/Fe, and/or $\bar{p}/p$. If reacceleration is included, the momentum-space diffusion coefficient $D_{pp}$ is related to the spatial coefficient $D_{xx}=\beta D_{0}\rho^{\delta}$ (Seo and Ptuskin 1994), where $\beta=v/c$ is the particle velocity, $\rho$ is the magnetic rigidity, and $\delta=1/3$ for a Kolmogorov spectrum of interstellar turbulence (Kolmogorov 1941), or $\delta=1/2$ for an Iroshnikov-Kraichnan cascade (Iroshnikov 1964; Kraichnan 1965), but can also be arbitrary. The spatial diffusion coefficient can also depend on position. This option was developed and utilised by Ackermann et al. (2015), where the positional dependence of the spatial diffusion coefficient was linked to the distribution of the Galactic magnetic field (GMF) strength.
@@
The GalProp code computes a complete network of primary, secondary, and tertiary isotope production starting from input CR source abundances. The nuclear reaction network is built using the Nuclear Data Sheets (2018), where a detailed description of its method of construction is given by Boschini et al. (2020, 2021a). Included are multistage chains of $p$-, $n$-, $d$-, $t$-, 3He-, $\alpha$-, and $\beta^{\pm}$-decays, and K-electron capture, as well as, in several cases, more complicated reactions. This accounts for up to 81 daughter nuclei in the final state for each fragment produced in the spallation of the target nucleus, plus an unlimited number of $p$-, $n$-, and $\beta^{\pm}$-decays.
Because the decay branching ratios and half-lives of fully stripped and hydrogen-like ions may differ (a well-known example is 7Be), GalProp includes the processes of K-electron capture, electron pickup from the neutral ISM gas, and formation of hydrogen-like ions as well as the inverse process of electron stripping (Pratt et al. 1973; Wilson 1978; Crawford 1979). Meanwhile, the fully stripped and hydrogen-like ions are treated as separate species.
The isotopic production cross section routines are built using LANL nuclear codes (Moskalenko et al. 2001; Moskalenko and Mashnik 2003; Mashnik et al. 2004; Moskalenko et al. 2005; Blann 1996; Blann and Chadwick 1998), databases (LANL, EXFOR [^2] https://www-nds.iaea.org/exfor/), an extensive literature search (Boschini et al. 2021a), and parameterizations (Webber et al. 2003; Silberberg et al. 1998; Tsao et al. 1998; Letaw et al. 1983). Cross sections for production of 2H, 3H, 3He, and total inelastic cross sections parameterizations (Wellisch and Axen 1996; Barashenkov and Polanski 1994) and corrected Tripathi formalism are described in Porter et al. (2022, Appendices F, G). The latest updates of the production cross sections are summarised in Cummings et al. (2025, Appendix B.4).
@@
For 2D calculations, analytical models for the gas density distribution are available (Moskalenko et al. 2002) (GP2D). The radial distribution for H i is taken from Gordon and Burton (1976) while the vertical distribution is from Dickey and Lockman (1990) for galactocentric radial distances $0\leq R\leq 8$ kpc and Cox et al. (1986) for $R\geq 10$ kpc with linear interpolation in between. The CO gas distribution is taken from Bronfman et al. (1988) for 1.5 kpc<<RR<<10 kpc, and from Wouterloot et al. (1990) for $R\geq$10 kpc, and is augmented with the Ferrière et al. (2007) model for $R\leq$1.5 kpc. Both of the 2D atomic and molecular gas distributions are rescaled to the common IAU recommended Sun-Galactic centre (GC) distance of $R_{\odot}=8.5$ kpc (Kerr and Lynden-Bell 1986). The H ii gas distribution is given by the NE2001 model (Cordes and Lazio 2002, 2003; Cordes 2004) with the updates from Gaensler et al. (2008).
For 3D simulations, the H i and 12CO distributions from Jóhannesson et al. (2018) (J18) are available. These were developed using a maximum-likelihood forward-folding optimisation applied to the LAB-H i (Kalberla et al. 2005) and CfA composite CO data (Dame et al. 2001; Dame and Thaddeus 2004). New with this paper we also include the gas distribution from Söding et al. (2025) (S25) who made a reconstruction of a very similar collection of data to the J18 work using 3D Gaussian processes to model correlations in the interstellar gas over different sight lines to enforce coherent structure for its distribution. Compared to the 2D models, the added degrees of freedom allow the optimised distributions to better reproduce the features observed in the line-emission surveys.
Multiple 2D and 3D models for the interstellar radiation field and GMF are also available with GalProp. However, these affect the propagation/energy losses for the CR electrons/positrons, which are not investigated in the current work. We therefore refer the interested reader to Porter et al. (2022) and associated references for more details.
III Calculations
III.1 Model Configurations
Figure 1: Model configuration components. Left panel: profiles along the $Y=0$ kpc direction for the CR source (solid line, arbitrary units) and GP2D gas surface densities (H i, dash-dotted line and H2, long dashed line). Right panel: S25 model gas surface density for the nearby Galactic neighbourhood. The Solar system is marked by the Sun symbol in the centre, and the dashed line is a circle of radius 200 pc about it.
@@ III.1 Model Configurations
The baseline model is 2D Galactocentric axisymmetric with a spatial model for the CR sources following the distribution of pulsars from Lorimer et al. (2006). The injection spectrum for the sources is a power-law in rigidity with a single injection index for all species below a break at 3 GV and a different common index for all species above that. The interstellar gas distribution is the standard 2D model GP2D described above. The propagation model uses a single power-law index $\delta$ for the diffusion coefficient $\propto D_{0}\rho^{\delta}$, where $D_{0}$ is the normalisation at 4 GV and $\rho$ is rigidity. The size of the Galaxy has maximum radius of 20 kpc with a $Z$-boundary (halo size) 4 kpc.
For the 3D models, the PD phenomenology and rigidity dependence for the diffusion coefficient is the same as the baseline. The size of the Galaxy has a maximum $X/Y$ boundary at $\pm$20 kpc and $Z$-boundary at 4 kpc. For the gas spatial distribution we use the GP2D and S25 models described above. The molecular gas density for the S25 model is obtained using the baseline/GP2D $X_{CO}$ distribution. The source density model is the same as the baseline. For the 3D solutions we also consider a spherical hole in the source density about the Solar system. With these models we test the effect of smooth and inhomogeneous source and gas distributions for the nearby Galactic neighbourhood.
@@
All model configuration files necessary to reproduce our results will be available as supplementary material from the GalProp website as is our usual practice.
III.2 CR Solution Geometry
Figure 2: Model combination results together with V1 data for 2013/1–2021/365 in the VLISM from Cummings et al. (2025) for elemental spectra Z=1–10. Line styles: 2D, 3D/0, solid; 3D/100, short dashed; 3D/200, long dashed; S25/0, dotted; S25/100, short dash dotted; S25/200, long dash dotted.
@@ III.2 CR Solution Geometry
For 3D models, we calculate the CR solution using a spatial grid with the non-uniform tan grid functions for all coordinates $X,Y,Z$, as described by Appendix E of Porter et al. (2022). The parameters for the individual coordinate functions are chosen so that the pixel size close to the Solar system is $\sim$12 pc. For the $X/Y$ coordinates the parameters are chosen so that the resolution slowly decreases with distance from the Solar system until it reaches the size of the $Z$ (halo) boundary, and then coarsens more beyond that to the respective planar boundaries. Because the gas distribution falls rapidly with increasing $Z$, the grid function parameters are chosen so that the resolution slowly decreases only until a distance 1 kpc above the plane and then coarsens more beyond that to the halo boundary.
@@
III.3 Tuning Procedure
The 2D baseline model was obtained by Cummings et al. (2025) using the GalFit package. This is a tool that iteratively optimises a GalProp model configuration parameter set against a user-specfied collection of data. For the 2D case this is computationally feasible because the likelihood evaluations are very quick, allowing for the parameter space to be adequately sampled. Such a procedure is not feasible for the high-resolution 3D models in this paper because the computational costs are much higher.
@@ III.3 Tuning Procedure
III.4 Results
Figure 3: Model combination results together with V1 data for 2013/1–2021/365 in the VLISM from Cummings et al. (2025) for B/C. Line styles as for Fig. 2.
Figure 4: Model combination results together with V1 data for 2013/1–2021/365 in the VLISM from Cummings et al. (2025) for elemental spectra Z=11–19 (left) and Z=20–28 (right). Line styles as for Fig. 2.
Figure 5: Elemental ratios for He, O, and Fe compared with V1 data. For each ratio, the corresponding V1 ratio has been derived from the spectra using power-law interpolation at the mid-bin energy for the elemental denominator species. Uncertainties are obtained by combining in quadrature the 1-sigma uncertainties for each elemental species interpolated using the same method. Left panel: ratios for combinations with 3D/GP2D and S25 gas models with source density hole sizes 0, 100, and 200 pc. Right panel: ratios for S25 gas model with source density hole sizes 0, 100, 150, 200, and 400 pc. Line styles are same as for Fig. 2, with the addition of double-dot dashed and triple-dotted dashed lines for 150 pc and 400 pc source density holes, respectively.
Figure 6: Model results for baseline and S25/150 with Boron pure secondary (4022a) and with primary Boron following 4022c abundances obtained by Cummings et al. (2025) together with V1 data for 2013/1–2021/365 in the VLISM. Panels: B/C, left; Be/O, right. The V1 Be/O ratio is derived using power-law interpolation at the geometric mean energy for the elemental Be spectrum. Line styles as for Fig. 5, with orange lines showing 4022c model curves.
Figure 2 shows the results for the elemental spectra for Z=1–10 with the V1 data from Cummings et al. (2025). The different line styles correspond to the GP2D and S25 gas models with different source density hole radial sizes: no hole, 100 pc, and 200 pc. It can be seen that the modelled CR spectra for all configurations closely agree for energies $\gtrsim$100–200 MeV nucleon-1 showing consistent normalisation. Differences occur for energies below this according to the gas model used and source density hole size.
Between the gas models, the CR intensities for the homogeneous (no-hole) source densities are different. While the GP2D/3D combination is as for the 2D baseline fit, for the S25/0 and S25/100 configurations the low-energy spectra are considerably higher. The CR injection occurs right up to the immediate vicinity of the Solar system when there is no hole. The inhomogeneous gas distribution is particularly low ($\lesssim$100 pc) nearby, and this results in slower energy losses. Consequently, the spectra are not as attenuated as for the homogeneous gas distribution with the GP2D model.